The total of three sisters' ages is 39. Dina is half as old as Michelle and 3 years younger than Juliette. How old are the sisters?

Answer by TutorDelphia(193)

Assign variables based on their names.
We know that:
d+m+j=39
We also know that
d=1/2m
and
d=j-3
We can find from that that
m=2d
and
j=d+3
Lets sub those in
d+2d+d+3=39
4d+3=39
4d=36
d=9
therefor j=12 and m=18